OpenStudy (anonymous):

You are home from school steadily at 95 km/h for 130 km .It then begins to rain and you slow down to 65km/h.You arrived home after driving 3hrs and 20 minutes. a)How far is your hometown from school? b)What was your average speed?

OpenStudy (anonymous):

My teacher got Finding :total distance speed What is given in the question is:distance =130km(d1) speed= s1=95km/h speed 2 =s2=65km/h t1-time 1=3.20minutes

OpenStudy (anonymous):

3 +20/60 =3.33s d1=130km d2=? solution: s2=d2/t2 initial t1=? second t2+?

OpenStudy (anonymous):

why did he put s2=d2/t2 i thought the question is asking for distance , since it says how far..

OpenStudy (abmon98):

Time driving 95km/h can be found with (speed=Distance/Time), (Speed*Time=Distance), (time = distance/speed) Time = 130 km/95km/hr = 1.37 hours or 1 hour 22 min. Total time of trip is 3 hr. 20 min. So, time driving 65 km/hr = 1 hour 58 min Distance traveled while driving 65km/hr can be found with (d = speed x time) Distance = 65 km/hr x 1.97 hours = 128 km Total distance traveled= 130 km + 128 km = 258 km Average Speed=Total Distance/Total Time A.Speed=258 km/3.33 hours